### v3

# Token-based Audio Inpainting via Discrete Diffusion

Abstract
Audio inpainting seeks to restore missing segments in degraded recordings. Previous diffusion-based methods exhibit impaired performance when the missing region is large. We introduce the first approach that applies discrete diffusion over tokenized music representations from a pre-trained audio tokenizer, enabling stable and semantically coherent restoration of long gaps. Our method further incorporates two training approaches: a derivative-based regularization loss that enforces smooth temporal dynamics, and a span-based absorbing transition that provides structured corruption during diffusion. Experiments on the MusicNet and MAESTRO datasets with gaps up to 750 ms show that our approach consistently outperforms strong baselines across range of gap lengths, for gaps of 150 ms and above. This work advances musical audio restoration and introduces new directions for discrete diffusion model training. Audio examples of our proposed method can be found at https://iftach21.github.io/.
11footnotetext: Equal contribution
1 Introduction
Audio inpainting refers to the task of reconstructing missing or corrupted segments of an audio signal (adler2011audio, ). It is a fundamental inverse problem in audio processing, with applications ranging from restoring damaged recordings and removing artifacts, to filling in gaps caused by data loss in transmission or editing operations (moliner2023diffusion, ). Traditional approaches to this problem often relied on signal modeling techniques such as autoregressive models (adler2011audio, ), sparse representations (marafioti2020gacela, ), or linear predictive coding (adler2011audio, ). While effective under certain assumptions (e.g., local stationarity), these methods typically perform well only on short gaps and may struggle with long-range dependencies or semantic coherence (moliner2023diffusion, ; zhou2019vision, ).
In recent years, deep generative models have significantly advanced the state of audio inpainting. Models such as Variational Autoencoders (VAEs) (marafioti2019context, ) and diffusion probabilistic models (kong2020diffwave, ; liu2023maid, ; ho2020denoising, ) have demonstrated the ability to learn expressive priors directly from large-scale audio datasets. Notably, diffusion models have emerged as particularly powerful tools for solving ill-posed inverse problems due to their iterative denoising process and strong generative capacity. Approaches such as DiffWave (kong2020diffwave, ) apply diffusion directly on waveform samples, while others such as MAID (liu2023maid, ) and CQT-Diff+ (moliner2023diffusion, ) operate in the continuous time-frequency domain using spectrograms or Constant-Q Transform (CQT) representations (moliner2023diffusion, ; zhou2019vision, ).
While deep generative approaches have achieved strong results, they each face challenges on long gaps. Waveform-level diffusion models (e.g., DiffWave) operate at the raw sample rate, making it difficult to capture long-range structure without very large receptive fields. Spectrogram and CQT-based methods (e.g., MAID, CQT-Diff+) use efficient time-frequency representations but depend on phase reconstruction and often lose coherence across extended regions. Similarly, VAE-based models preserve local continuity but struggle to maintain semantic consistency as the gap length increases. These limitations highlight the difficulty of reconciling fine temporal detail with high-level structure in continuous audio representations.
In this work, we propose a novel approach: Audio Inpainting via Discrete Diffusion (AIDD). Unlike prior methods that operate in the continuous domain (kong2020diffwave, ; liu2023maid, ; moliner2023diffusion, ; zhou2019vision, ), our method applies the diffusion process to a discrete latent space. Specifically, we employ the WavTokenizer (ji2024wavtokenizer, ) to quantize audio signals into compact sequences of discrete tokens, and perform the diffusion process entirely in this categorical space. This formulation allows the model to capture high-level semantic structures in audio, while avoiding the challenges of modeling raw waveforms or spectrograms directly. To address these challenges, we introduce a discrete diffusion framework for audio inpainting. Our contributions are threefold, advancing audio inpainting with discrete diffusion model:
  • Discrete diffusion for music. To the best of our knowledge, our proposed method AIDD, is the first to apply discrete diffusion to tokenized music representations, leveraging the WavTokenizer to transform raw waveforms into compact token sequences. This enables stable generation and long-range semantic coherence.
  • Span-based masking with derivative-regularized reconstruction loss. We introduce a unified corruption-reconstruction framework that combines structured span masking with a smoothness-oriented regularization loss. The forward process masks contiguous spans of tokens, enabling a progression from fine-grained local corruption to broader semantic perturbations. Additionally, during reconstruction, a derivative-based loss encourages continuity across tokens, penalizing irregular local variations and yielding more natural and perceptually faithful inpainting.
  • Performance. We present comparable or state-of-the-art results for audio inpainting on MusicNet (dataset, ) and MAESTRO (MAESTRO, ) dataset, for gaps ranging from 150 - 750 ms. Showing consistent results for both large and small datasets.
2 Related Work
2.1 Methods for Audio Inpainting
The term “audio inpainting” was introduced in (adler2011audio, ) to describe the restoration of missing or corrupted audio segments, although related notions had appeared under names such as audio interpolation (marafioti2020gacela, ), audio extrapolation (lieb2018audio, ), missing sample reconstruction, waveform substitution (adler2011audio, ), and imputation (lieb2018audio, ).
Classical methods mainly targeted short gaps (<100<100 ms), often assuming local stationarity. Auto-regressive models (adler2011audio, ) predicted missing samples from past ones but degrade on longer gaps. Sparse reconstruction in time–frequency domains like STFT or Gabor (lieb2018audio, ; adler2011audio, ) exploited algorithms such as OMP (tropp2007signal, ), later improved by adaptive dictionaries (e.g., learned Gabor atoms (taubock2020dictionary, )). NMF-based methods (mokry2023algorithms, ) filled spectrogram gaps via low-rank factorization, with probabilistic variants improving robustness. Other structured priors include sinusoidal modeling for harmonic signals (adler2011audio, ) and graph-based regularization: (perraudin2018inpainting, ) used a graph Laplacian on self-similarity matrices to borrow content from uncorrupted regions.
While effective for very short gaps, these models fail for longer ones due to violated assumptions of stationarity or sparsity. Deep learning approaches were introduced to overcome this challenge. CNNs first inpainted spectrograms, with (marafioti2019context, ) using a U-Net style context encoder for tens-of-milliseconds gaps. GANs further improved realism: (ebner2020audio, ) proposed a Wasserstein GAN with multi-context conditioning for gaps up to 500 ms, while (marafioti2020gacela, ) extended this with GACELA, a multi-scale GAN handling 1–1.5 s gaps. These data-driven models substantially outperform classical approaches, especially on long gaps with complex temporal or musical structure. More recently, A score-aware GAN framework (aironi2023score, ) further improved music inpainting by incorporating symbolic losses, outperforming GACELA on medium gaps.
2.2 Audio Inpainting using Diffusion Models
Diffusion models, or score-based generative models, have recently transformed vision and audio generation. They define a forward noising process and learn its reversal (e.g., DDPMs (ho2020denoising, ), score-SDEs (song2020score, )). In audio, they support waveform and spectrogram generation conditioned on context. For instance, DiffWave (kong2020diffwave, ) synthesizes raw waveforms for vocoding, while other works apply diffusion in latent or time–frequency domains (STFT, CQT) to exploit structure (moliner2023solving, ; moliner2023diffusion, ). Most use continuous diffusion, though discrete approaches exist: DiffSound (yang2023diffsound, ) applies discrete diffusion to quantized spectrogram tokens. Conceptually, continuous diffusion perturbs real values, yielding smooth interpolations, whereas discrete diffusion masks/replaces categorical tokens, enabling multimodal reconstructions but with possible quantization artifacts (ho2020denoising, ; song2020score, ; yang2023diffsound, ).
Diffusion has recently been adapted to inpainting. (moliner2023diffusion, ) propose a zero-shot unconditional model on CQT spectrograms, with posterior sampling to enforce data consistency (kawar2022denoising, ; chung2022diffusion, ). This handles arbitrary gaps, matching baselines on short ones and surpassing them up to 300 ms. Similarly, CQT-Diff (moliner2023solving, ) uses an invertible CQT front-end for inpainting, declipping, and bandwidth extension, achieving state-of-the-art restoration. MAID (liu2023maid, ), instead, trains a conditional diffusion model for music inpainting guided by context and auxiliary cues (e.g., pitch), enabling longer gaps but requiring task-specific supervision. (aironi2023score, ) further propose score-aware conditioning, aligning the score function with masked regions for musically coherent results.
Architectures differ in operating domain (raw waveform vs. spectrogram/latent space) and conditioning. Zero-shot approaches (moliner2023diffusion, ) enforce context consistency during inference, while conditional models integrate masked audio or side information during training. Masking usually involves contiguous spans: Moliner et al. test gaps from 25–300 ms, while GAN baselines address seconds-long gaps (marafioti2020gacela, ; ebner2020audio, ). Diffusion models degrade gracefully with gap length: all interpolate short gaps well, but diffusion’s generative prior yields more plausible completions beyond 100 ms compared to sparsity- or AR-based methods.
In summary, diffusion-based inpainting combines strong generative priors with flexible conditioning. Open questions remain on continuous vs. discrete diffusion (yang2023diffsound, ) and conditioning strategies (posterior sampling vs. trained conditional networks).
3 Preliminary
3.1 Discrete Diffusion Models
Continuous Diffusion Models (CDMs) (ho2020denoising, ; song2020score, ) have achieved remarkable success across various generative modeling tasks, particularly in the image domain (dhariwal2021diffusion, ). The high-dimensional nature of raw audio data adds to the difficulty of directly modeling it within the diffusion framework. An effective strategy to address this issue is to first compress continuous audio signals into compact, discrete representations, such as quantized tokens derived from a learned codebook (defossez2022high, ). DDMs, which operate in a token space, have recently demonstrated strong performance in the domain of natural language generation (lou2310discrete, ; nie2025large, ; sahoo2024simple, ). Similarly to CDM, DDMs comprise a forward noising process that progressively corrupts an initial sample $x_{0}$ into a maximally noisy (fully masked) version $x_{T}$, and a reverse denoising process trained to reconstruct $x_{0}$ from $x_{T}$.
Forward process.
Given the tokenized audio data $x_{0}\sim p_{\text{data}}$ over a finite alphabet $\mathcal{X}=\{1,\ldots,N\}$, the forward process is defined by a transition matrix $Q_{t}$ that indicates how $x_{t-1}$ transits to $x_{t}$ for each step in the forward process. The behavior of the process is given by:
$\mathbb{P}(x_{t+\Delta t}=y\mid x_{t}=x)=\delta_{xy}+Q_{t}(y,x)\Delta t+\mathcal{O}(\Delta t^{2}),$ (1)
where $Q_{t}(y,x)$ denotes the rate of transitioning from state $x$ to state $y$ at time $t$.
Existing work (lou2310discrete, ) shows that the forward diffusion in discrete space can be formulated using a token-level transition rate matrix $Q_{\text{tok}}\in\mathbb{R}^{n\times n}$, where $n$ is the vocabulary size. The forward transition probability of a single token $x_{0}$ at time $t$ is then given by
$p^{\text{tok}}_{t|0}(\cdot\mid x_{0})\;=\;\exp\!\big(\bar{\sigma}(t)\,Q_{\text{tok}}\big),$ (2)
where $\bar{\sigma}(t)=\int_{0}^{t}\sigma(s)\,ds$ denotes the cumulative noise and $Q_{\text{tok}}\in\mathbb{R}^{n\times n}$ is a fixed matrix of absorbing design, where all tokens eventually transition into a terminal [MASK] token.
When $Q_{\text{tok}}$ is set to the absorbing mask transition matrix, the process reduces to only two possible outcomes: the token stays unchanged with probability $e^{-\bar{\sigma}(t)}$, or it is replaced by the [MASK] token with probability $1-e^{-\bar{\sigma}(t)}$.
Reverse Process.
The reverse process is described by a new reverse transition matrix (sun2022score, ; kelly1981reversibility, ) which is defined as $\bar{Q}_{t}(y,x)=\tfrac{p_{t}(y)}{p_{t}(x)}Q_{t}(x,y)$ and $\bar{Q}_{t}(x,x)=-\sum_{y\neq x}\bar{Q}_{t}(y,x)$.
To simulate the reverse process, it is sufficient to estimate the ratio $\frac{p_{t}(y)}{p_{t}(x)}$, referred to as the concrete-score (lou2310discrete, ; meng2022concrete, ), which can be approximated by a neural score function: $s_{\theta}(x,t)\approx\left[\tfrac{p_{t}(y)}{p_{t}(x)}\right]_{y\in\mathcal{X},\,y\neq x}$.
Training.
To train this score function, recent work has proposed the Diffusion Weighted Denoising Score Entropy (DWDSE) objective (lou2310discrete, ), which is defined as:
$\mathcal{L}_{\text{DWDSE}}=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot\mid x_{0})}\left[\sum_{y\neq x_{t}}Q_{t}(x_{t},y)\left(s_{\theta}(x_{t},t)_{y}-\frac{p_{t|0}(y\mid x_{0})}{p_{t|0}(x_{t}\mid x_{0})}\log s_{\theta}(x_{t},t)_{y}+C\right)\right]dt,$ (3)
where $C=K\left(\frac{p_{t|0}(y\mid x_{0})}{p_{t|0}(x_{t}\mid x_{0})}\right)$, and $K(a):=a\log a-a$ is a convex regularization term. Once trained, we can construct $\mathbf{x}_{t-\Delta t}$ by sampling each token $x_{t-\Delta t}^{i}$ from the probability
$p(x_{t-\Delta t}^{i}\mid x_{t}^{i})\;=\;\delta_{x_{t}^{i}}\!\left(x_{t-\Delta t}^{i}\right)+\Delta t\,Q_{t}^{\text{tok}}\!\left(x_{t}^{i},x_{t-\Delta t}^{i}\right)s_{\theta}(\mathbf{x}_{t},t)_{i,\,x_{t-\Delta t}^{i}}.$ (4)
In this way, by iteratively applying the reverse step, the model is able to generate new samples.
4 Audio Inpainting using Discrete Diffusion Model
To the best of our knowledge, our method Audio Inpainting using Discrete Diffusion (AIDD) is the first to leverage a discrete diffusion model for audio inpainting. AIDD comprises three main components: (1) audio tokenization using WavTokenizer (ji2024wavtokenizer, ), (2) generative modeling using a DDM (lou2310discrete, ), and (3) waveform reconstruction via token decoding (ji2024wavtokenizer, ). Furthermore, we propose a new training derivative-based loss, and a span-based masking method, extending the approach of (lou2310discrete, ) to more effectively address the audio inpainting task.
4.1 The Proposed Architecture
Method Overview.
AIDD consists of WavTokenizer (ji2024wavtokenizer, ) that transforms high-dimensional raw audio signals into compact sequences of discrete tokens and reconstructs it back to raw audio, and a Diffusion Transformer (DiT) architecture (peebles2023scalable, ) which learns to predict masked tokens through the reverse diffusion process. An overview of the proposed framework is illustrated in Figure 1.
Figure 1: Our method operates on audio signals with missing (silent) segments. During inference, the input waveform, containing a silence gap is processed by the WavTokenizer encoder, which converts the audio into a discrete sequence of tokens. Next, a DiT performs inpainting by iteratively predicting the masked tokens, resulting in reconstructed token sequence. Finally, the reconstructed tokens are passed through the WavTokenizer’s decoder to synthesize the output audio waveform in the masked part. During training, token sequences are corrupted with span-based masking at randomly sampled timesteps, and the DiT is optimized to predict the concrete score using the DWDSE objective, complemented by the derivative-based loss.
Audio Tokenization.
The first step in our method is to convert the raw audio waveform into discrete audio tokens that can be efficiently processed and trained by the DDM. We utilize a pre-trained audio tokenizer WavTokenizer (ji2024wavtokenizer, ), which compresses high-resolution audio into a compact sequence of discrete tokens using a single quantizer. Despite its extreme compression, WavTokenizer preserves high reconstruction fidelity and rich semantic content. Unlike conventional audio inpainting frameworks that operate directly on the audio waveform or the spectrogram (moliner2023diffusion, ), audio tokenization allows us to formulate audio inpainting as a discrete sequence completion task.
Discrete Diffusion Model.
At the core of our method is a DiT architecture (peebles2023scalable, ), which integrates time conditioning into a standard encoder-only transformer (vaswani2017attention, ; devlin2019bert, ) and, following the work of (lou2310discrete, ), we incorporate rotary positional encoding (su2024roformer, ). To operate effectively in the discrete token space, we use the denoising score entropy formulation, guiding the reverse diffusion dynamics as seen in Eq. 3.
4.2 Derivative-Based Regularization Loss for Smooth Token Predictions
While the DWDSE objective in Eq. 3 ensures that the score function $s_{\theta}$ learns correct transition ratios for masked tokens, it does not explicitly constrain the temporal consistency of predicted token embeddings across neighboring positions. To address this, we introduce a derivative-based regularization term that directly extends the formulation of Section 3 by imposing smoothness on token trajectories in embedding space.
Let $e_{i}\in\mathbb{R}^{d}$ denote the ground-truth embedding of the token at position $i$, and $\hat{e}_{i}\in\mathbb{R}^{d}$ its predicted embedding. We define discrete temporal derivatives for both ground-truth and predicted embeddings. For first-order differences:
$\Delta^{1}e_{i}=e_{i+1}-e_{i},\quad\Delta^{1}\hat{e}_{i}=\hat{e}_{i+1}-\hat{e}_{i},$
and for second-order differences, capturing local curvature:
$\Delta^{2}e_{i}=e_{i+1}-2e_{i}+e_{i-1},\quad\Delta^{2}\hat{e}_{i}=\hat{e}_{i+1}-2\hat{e}_{i}+\hat{e}_{i-1}.$
We then define a regularization loss that aligns the derivatives of predicted and ground-truth embeddings at masked locations:
$\mathcal{L}_{\text{deriv}}=\frac{1}{|M|}\sum_{i\in M}\big\|\Delta^{k}\hat{e}_{i}-\Delta^{k}e_{i}\big\|^{2},$
where $k\in\{1,2\}$ selects first- or second-order differences, and $M$ is the set of indices involving masked tokens (or their neighbors).
This additional term complements the DWDSE objective by penalizing irregular local fluctuations in predicted embeddings, thereby encouraging the reverse process to reconstruct sequences that respect the inherent smoothness of natural audio token trajectories. The overall training objective thus combines both components:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DWDSE}}+\lambda\,\mathcal{L}_{\text{deriv}},$
where $\lambda>0$ is a weighting factor that balances the contribution of the derivative-based regularization.
4.3 Span Masking
Whereas prior discrete diffusion frameworks (lou2310discrete, ) apply masking independently in the forward process, we propose a structured corruption strategy that enables the model to better handle inpainting. We employ a span-based masking strategy for the absorbing diffusion process.
Given the sequence of tokens $X=(x_{1},\dots,x_{L})$ of length $L$, we construct a subset of perturbed tokens by iteratively sampling contiguous spans until a predefined budget is met. The budget corresponds to the expected number of tokens transitioned away from their original state, defined as $B(t)=\big(1-e^{-\sigma(t)}\big)\cdot L$, where $\sigma(t)$ denotes the noise schedule at time $t$, and $1-e^{-\sigma(t)}$ represents the probability of transitioning away from the current state. This formulation ensures that the proportion of perturbed tokens aligns with the transition probability induced by the forward diffusion process.
At each iteration, we sample a span length $\ell\sim\mathrm{Geo}(p_{\sigma})$, where $p_{\sigma}$ controls the bias toward shorter spans and is defined as $p_{\sigma}=\tfrac{p_{0}}{1+\alpha\sigma}$, with base parameter $p_{0}$ and scaling factor $\alpha$ determining how the noise level $\sigma$ influences the expected span length. Thus, early timesteps favor short spans, with longer spans becoming more likely later, capped by $\ell_{\max}$. For each sampled span, a starting index is drawn uniformly from $1,L-\ell]$, and all tokens in the span are masked.
5 Experiments
5.1 Objective Metrics
We adopt a combination of metrics for evaluations, capturing both the technical fidelity and perceptual plausibility of the reconstructed audio.
Fréchet Audio Distance (FAD).
We use FAD (kilgour2018fr, ), which measures the distributional distance between real and generated audio features. It reflects perceptual similarity in attributes such as timbre and texture, serving as an indicator of realism. Our evaluation uses a VGG-based feature extractor.
Objective Difference Grade (ODG).
Perceptual quality is further assessed with ODG from the PEMO-Q model (huber2006pemo, ), which simulates auditory processing and maps a Perceptual Similarity Measure (PSM) to scores from 0 (imperceptible) to –4 (very annoying). We also assessed with ODG using PEA-Q. Both were included to match the evaluation methods of related works. As a reference-based metric aligned with listening tests, ODG complements waveform-based measures.
Log Spectral Distance (LSD).
Following (moliner2023diffusion, ), we compute LSD to capture spectral distortions:
$\text{LSD}=\frac{1}{T}\sum_{t=1}^{T}\sqrt{\frac{1}{K}\sum_{k=1}^{K}\left(\log|X_{t,k}|^{2}-\log|\hat{X}_{t,k}|^{2}\right)^{2}}$ (5)
where $X_{t,k}$ and $\hat{X}_{t,k}$ are the STFTs of the original and reconstructed signals. We use a window of $K=2048$ samples and hop length 512.
5.2 Experimental setup
Dataset.
We evaluated our model on two benchmark datasets. First, we used the MusicNet dataset (dataset, ), which provides a predefined split of training and test sets. The training set comprises 330 freely licensed classical music recordings with aligned labels for instruments, notes, and other musical attributes. This dataset serves as a compact benchmark for evaluating models on polyphonic music transcription. We report results using the predefined split.
In addition, we evaluated our model on the MAESTRO dataset (MAESTRO, ), which also provides predefined training and test splits. MAESTRO contains over 200 hours of virtuosic piano performances with precise note-level alignment between audio and MIDI. Due to its large size, we trained only on a subset of the official training set (see Supplementary Material for details).
We conducted experiments on each dataset independently; there was no mixing or combined training across datasets.
Training.
During training, raw audio signals are first tokenized and truncated to a fixed length of 1024 tokens (approximately 13.6 seconds). A random timestep is selected, and noise (in the form of token masking) is applied corresponding to the span-based masking corruption. The DDM is then trained to predict the concrete scores from this corrupted input. Through reverse diffusion steps, the model gradually denoises the token sequence, effectively learning to recover clean representations from noisy ones in token space, illustrated in Figure 1.
Training employed the AdamW optimizer (learning rate of $10^{-6}$), batch size 128, and 1024 token samples. On MusicNet, the base model with DWDSE loss was trained for 400k steps (two days) on a single NVIDIA A6000 GPU, while other methods were trained for 100k steps. On MAESTRO, training ran for 150k steps (24h) under the same setup. Full hyperparameters are reported in the supplementary, following ablation results (see Section 7).
Evaluation.
We evaluate the effectiveness of our proposed audio inpainting method, we followed the experimental protocol of (moliner2023diffusion, ) and selected 60 previously unseen music segments from the MusicNet test set (dataset, ), each with a duration of 4.17 seconds. For each segment, we introduced four synthetic gaps at fixed, evenly spaced locations. The gap durations varied across experiments, ranging from 150 ms to 300 ms. Since AIDD is stochastic, we generated 10 samples for each segment and reported the averaged score as the final result.
For the MAESTRO dataset (MAESTRO, ), we followed (aironi2023time, )’s experimental set-up, taking 6 seconds segments from the predefined test set and introducing single gaps in the middle.
These corrupted inputs were then processed using our method AIDD (see Section 5.2), which reconstructed the missing audio to produce complete, inpainted waveforms. We evaluated the reconstructed samples against the corresponding ground truths using the objective metrics described in Section 5.1.
Inference Pipeline.
At inference, the masked waveform is tokenized with the pretrained WavTokenizer encoder. The trained DDM then performs reverse diffusion to predict missing tokens conditioned on context. The token sequence is decoded back to waveform, and only the inpainted segment replaces the original gap, avoiding unnecessary re-encoding of intact regions. A short 6 ms crossfade is applied at boundaries to ensure smooth transitions. An overview is shown in Figure 1.
6 Results
For the MusicNet dataset, we evaluate our method against three baselines: LPC, an extrapolation method using linear predictive coding (kauppinen2002audio, ); A-SPAIN-L, a sparsity-based inpainting approach with dictionary learning (taubock2020dictionary, ); and CQT-Diff+, a diffusion model with Constant-Q Transform, state-of-the-art for gaps up to 300 ms (moliner2023diffusion, ).
For the MAESTRO dataset (MAESTRO, ) we evaluated against three baselines as well: GACELA (marafioti2020gacela, ): a cGAN that performs long-gap audio inpainting using multiple discriminators at different time scales; bin2bin (aironi2023score, ): a spectrogram inpainting model based on a pix2pix-style GAN with STFT losses; bin2bin-MIDI (aironi2023time, ): an extension of bin2bin that adds a piano-roll (MIDI) loss to enforce pitch consistency.
Table 1 reports the average performance across varying gap durations on the MusicNet (dataset, ) dataset, while Table 2 presents the corresponding results on the MAESTRO (MAESTRO, ) dataset.
Method 150 ms 200 ms 250 ms 300 ms
FAD $\downarrow$ LSD $\downarrow$ ODG $\uparrow$ FAD $\downarrow$ LSD $\downarrow$ ODG $\uparrow$ FAD $\downarrow$ LSD $\downarrow$ ODG $\uparrow$ FAD $\downarrow$ LSD $\downarrow$ ODG $\uparrow$
Masked 16.001 0.555 -3.873 18.244 0.763 -3.881 23.583 0.971 -3.891 33.342 1.162 -3.897
LPC 3.172 0.184 -3.351 4.883 0.258 -3.467 7.934 0.336 -3.512 11.907 0.415 -3.550
A-SPAIN-L 6.121 0.198 -3.668 12.038 0.311 -3.767 16.181 0.445 -3.801 21.574 0.610 -3.818
CQT-Diff+ 1.525 0.164 -3.559 2.619 0.218 -3.651 3.202 0.272 -3.891 4.652 0.324 -3.711
AIDD 1.866 0.162 -3.215 2.391 0.209 -3.250 2.438 0.260 -3.274 3.549 0.297 -3.284
Table 1: Performance comparison of methods across different gap lengths using FAD, LSD, and ODG metrics, on MusicNet dataset (dataset, ). Lower is better for FAD/LSD ($\downarrow$), higher is better for ODG ($\uparrow$). In bold the best score. The AIDD presented is the highest reported scores, as described in section 5.2.
Method 375 ms ($\uparrow$) 750 ms ($\uparrow$)
GACELA -3.232 $\pm$ 0.232 -3.318 $\pm$ 0.202
bin2bin -2.892 $\pm$ 0.510 -3.039 $\pm$ 0.495
bin2bin-MIDI -2.800 $\pm$ 0.491 -2.976 $\pm$ 0.456
AIDD -2.370 $\pm$ 0.705 -2.959 $\pm$ 0.604
Table 2: ODG (PEA-Q) score values for the three compared methods to AIDD, on MAESTRO (MAESTRO, ) dataset. In bold the best score. Higher is better ($\uparrow$). The AIDD’s result reported is the combined method.
On the MusicNet dataset, our method (AIDD) consistently outperforms prior approaches across most gap durations. In particular, AIDD achieves the lowest FAD values for gaps of 200-300 ms, while also obtaining the best LSD and generally stronger ODG scores. Compared to the strong diffusion-based baseline CQT-Diff+, our approach yields significantly lower distortion (FAD/LSD) for medium and longer gaps, with improvements of nearly 25% in FAD at 300 ms (3.549 vs. 4.652). This suggests that AIDD generates more perceptually consistent reconstructions, especially as the gap size increases. At short gaps (150 ms), CQT-Diff+ achieves a lower FAD, though the margin is small; nevertheless, AIDD provides better ODG and LSD in this setting. Finally, it is worth noting that our base model is half the size of CQT-Diff+ and required only one-quarter of the training time, underscoring its efficiency.
On the MAESTRO dataset, results follow a similar trend (Table 2). Our method (AIDD) achieves the best ODG score for the 375 ms gap, outperforming all three baselines by a clear margin (2.370-2.370 vs. 2.800-2.800 to 3.232-3.232). This indicates that AIDD can better preserve perceptual quality even in the presence of interruptions. At 750 ms, our model also achieves the top performance (2.959-2.959), surpassing all three baselines, which was specifically designed for long-gap inpainting.
It is important to highlight that, unlike prior works which trained on the full MAESTRO dataset, we only used a reduced training subset, yet still achieved state-of-the-art performance for both shorter (375 ms) and longer (750 ms) gaps. This suggests that AIDD generalizes well to long-context musical structures while requiring fewer resources, both in terms of dataset size and computational budget.
Taken together, these findings confirm that AIDD is robust across datasets and gap durations: it provides high perceptual quality for medium and long gaps in MusicNet, and achieves strong performance on challenging, real-world piano recordings in MAESTRO.
7 Ablation Study
To further assess the effectiveness of AIDD, we conducted an ablation study on the MusicNet dataset, following the same experimental protocol described in Section 5.2. Specifically, we examined the impact of different training strategies: Derivative-Based (Section 4.2), Span-Based (Section 4.3), and their combination. All experiments were performed with the same architecture and inpainting setup to ensure fair comparison. The results are summarized in Table 3.
In addition, we investigated the effect of hyperparameter choices for the proposed training methods. Table 3 reports the performance of each variant, including the Derivative-Based, Span-Based, and the combined approach.
Overall, the combined method consistently yielded the best performance, achieving the highest scores across most evaluation settings. These results highlight both the robustness of the approach and its alignment with established audio inpainting evaluation protocols.
Method / Setting 200 ms 300 ms
FAD $\downarrow$ LSD $\downarrow$ ODG $\uparrow$ FAD $\downarrow$ LSD $\downarrow$ ODG $\uparrow$
AIDD (Base - DWDSE loss) 2.802 0.211 $\pm$0.05 -3.262 $\pm$0.06 4.015 0.303 $\pm$ 0.06 -3.296 $\pm$0.04
AIDD - Span-Based Variants ($\ell_{\max}=30$)
$p_{0}=0.6;\alpha=0.5$ 2.438 0.206 $\pm$0.05 -3.249 $\pm$0.06 3.573 0.292 $\pm$0.06 -3.284 $\pm$0.04
$p_{0}=0.8;\alpha=0.3$ 2.719 0.211 $\pm$0.05 -3.258 $\pm$0.06 3.626 0.298 $\pm$0.07 -3.289 $\pm$0.04
AIDD - Derivative-Based loss
$\lambda=200;\Delta^{1}e$ 2.455 0.209 $\pm$0.05 -3.251 $\pm$0.06 3.439 0.293 $\pm$0.06 -3.284 $\pm$0.05
$\lambda=500;\Delta^{1}e$ 2.705 0.211 $\pm$0.05 -3.255 $\pm$0.06 3.654 0.298 $\pm$0.07 -3.285 $\pm$0.04
$\lambda=800;\Delta^{1}e$ 2.869 0.215 $\pm$0.05 -3.260 $\pm$0.06 3.837 0.303 $\pm$0.07 -3.289 $\pm$0.04
$\lambda=200;\Delta^{2}e$ 2.593 0.210 $\pm$0.05 -3.251 $\pm$0.06 3.549 0.296 $\pm$0.07 -3.281 $\pm$0.05
$\lambda=500;\Delta^{2}e$ 2.599 0.213 $\pm$0.05 -3.255 $\pm$0.06 3.644 0.298 $\pm$0.07 -3.283 $\pm$0.04
$\lambda=800;\Delta^{2}e$ 2.539 0.213 $\pm$0.05 -3.256 $\pm$0.06 3.538 0.301 $\pm$0.07 -3.283 $\pm$0.05
AIDD - Combined methods (Derivative And Span Based) ($\ell_{\max}=30$)
$\lambda=500$ ; $p_{0}=0.8;\alpha=0.5;\Delta^{1}e$ 2.391 0.209 $\pm$0.05 -3.250 $\pm$0.06 3.549 0.297 $\pm$0.06 -3.284 $\pm$0.04
$\lambda=200$ ; $p_{0}=0.8;\alpha=0.5;\Delta^{1}e$ 2.518 0.209 $\pm$0.05 -3.254 $\pm$0.06 3.558 0.295 $\pm$0.07 -3.285 $\pm$0.05
$\lambda=500$ ; $p_{0}=0.8;\alpha=0.5;\Delta^{2}e$ 2.449 0.212 $\pm$0.05 -3.255 $\pm$0.06 3.488 0.302 $\pm$0.07 -3.287 $\pm$0.04
Table 3: Performance comparison of different AIDD methods across different gap lengths using FAD, LSD, and ODG metrics, on MusicNet dataset. In bold the best scores. Lower is better for FAD/LSD ($\downarrow$), higher is better for ODG ($\uparrow$).
8 Conclusion
In this work, we proposed a novel approach for audio inpainting by combining a DDM with a WavTokenizer-based encoder-decoder architecture (ji2024wavtokenizer, ). Rather than applying diffusion directly to raw waveforms, which are noisy and difficult to model, we operate in a discrete token space, enabling the DDM to effectively learn the underlying distribution of musical structure. This allows for more stable training and more coherent reconstruction during inference. Our method not only advances the task of audio inpainting, but also lays the foundation for additional generative capabilities, including unconditional music generation and music continuation based on a given context. Moreover, we introduce two novel training methods for the discrete diffusion process: span-based masking and derivative-based loss that better contribute to the model’s learning of both fine-grained and structured corruptions.
Experimentally, we demonstrated that our method achieves competitive or superior results compared to existing state-of-the-art approaches, particularly in challenging settings with longer gaps. While many previous methods focus on short gaps of up to 100 ms, our model maintains strong performance even on 750 ms gaps, highlighting its potential applicability to real-world scenarios where such gaps can significantly affect the listening experience. Overall, our method introduces a robust and extensible paradigm for music-based generative modeling in the token domain, with promising applications in both restoration and creative audio generation tasks, as well as lays the foundation for continuous work in the field of DDM training.
While our study focused on music inpainting, the proposed architecture is not limited to this domain. Future research will be extending the framework to speech audio inpainting, where missing segments can disrupt the intelligibility of entire sentences. Addressing such scenarios would demonstrate the broader applicability of discrete diffusion for real-world communication, and position this line of work as a practical solution for diverse audio restoration tasks.
Future work may extend our training strategies to large language models by building on our base method. Although audio and language differ inherently, both can be approached through token-based discrete diffusion.
Supplementary Material
Appendix A Code for AIDD
Our full code will be published in GitHub for open access once the paper is accepted. Currently, the code is added as a folder with the submission.
Appendix B Reproducibility Statement
With the following hyperparameter as well as appropriate hardware presented in our supplementary material, anyone can reproducibly conduct our experiment using our code which, upon acceptance, will be publicly accessed.
Appendix C Reproducibility Materials
C.1 Hyperparameters
Parameter Value
Learning rate 1e-6
Batch size 128
Sampling rate 128
Optimizer AdamW
$\lambda$ 500
$p_{0}$ 0.8
$\alpha$ 0.5
Derivative $\Delta^{1}e$
Table 4: Hyperparameters used in experiments.
C.2 Training Environment
We used singular NVIDIA A6000 GPU for all our experiments. The highest score presented was trained for just less than one day, for 100,000 steps for MusicNet dataset and for the MAESTRO dataset, it was trained for a day, with 150,000 steps.
Appendix D MAESTRO dataset
Due to its large size, we randomly took a subset of the training and test, following the test-train split predefined. Upon request, we will provide the full list.
Appendix E ODG Implementations: PEMO-Q vs. PEA-Q
In this work we employed two variants of the Objective Difference Grade (ODG): PEMO-Q and PEA-Q. While both provide perceptual quality estimates on the standardized scale from 0 (imperceptible) to -4 (very annoying), they are based on different models and implementations.
PEMO-Q.
PEMO-Q [11] is based on a detailed computational model of the auditory system, simulating peripheral and central auditory processing. The ODG values are obtained by mapping a Perceptual Similarity Measure (PSM) to the ITU-defined scale.
PEA-Q.
PEA-Q follows the ITU-R Recommendation BS.1387 (1999) [13] for objective audio quality assessment, which was designed as a standardized benchmark for audio codec evaluation. In our experiments, we used the MATLAB implementation.
These implementations allowed us to ensure consistency with prior work and comparability across different evaluation protocols.